# From Dataset to Real-world: General 3D Object Detection via Generalized Cross-domain Few-shot Learning

### v1

@@ Abstract
LiDAR-based 3D object detection datasets have been pivotal for autonomous driving, yet they cover a limited range of objects, restricting the model’s generalization across diverse deployment environments. To address this, we introduce the first generalized cross-domain few-shot (GCFS) task in 3D object detection, which focuses on adapting a source-pretrained model for high performance on both common and novel classes in a target domain with few-shot samples. Our solution integrates multi-modal fusion and contrastive-enhanced prototype learning within one framework, holistically overcoming challenges related to data scarcity and domain adaptation in the GCFS setting. The multi-modal fusion module utilizes 2D vision-language models to extract rich, open-set semantic knowledge. To address biases in point distributions across varying structural complexities, we particularly introduce a physically-aware box searching strategy that leverages laser imaging principles to generate high-quality 3D box proposals from 2D insights, enhancing object recall. To effectively capture domain-specific representations for each class from limited target data, we further propose a contrastive-enhanced prototype learning, which strengthens the model’s adaptability. We evaluate our approach with three GCFS benchmark settings, and extensive experiments demonstrate the effectiveness of our solution for GCFS tasks. The code will be publicly available.
@@ 1 Introduction
LiDAR-based 3D object detection [29, 23] has made considerable progress in recent years, largely driven by the availability of open-source datasets collected from various regions worldwide [10, 1, 36, 11]. These datasets have played a crucial role in advancing autonomous driving by providing foundational data for model training and evaluation. However, their coverage remains limited to a narrow range of object categories, often lacking the diversity needed to support generalization across various geographic and cultural contexts. Table 1 provides an overview of these datasets, highlighting the differences in object-detection focus. As models trained on these datasets are deployed in different regions or countries, they may encounter novel, region-specific objects that were not present in the original training data. Examples include electric scooters in China, tuk-tuks in Thailand, and indigenous wildlife in Africa.
In real-world applications of autonomous driving, new environments and novel object categories constantly emerge. Relying on large-scale datasets for each new environment to retrain or fine-tune a model from scratch is both time-consuming and resource-intensive, making it impractical for rapid deployment needs. Existing approaches on LiDAR-based 3D object detection, such as semi-supervised learning [41] and 3D open-vocabulary detection (OVD) [6, 51, 2], assume the availability of large amounts of unlabeled target-domain data, which may not always be feasible in deployment scenarios. Furthermore, while 3D-domain adaptation (DA)  [40, 48, 13] focuses on addressing domain shifts, it does not explicitly account for novel object categories that were unseen during training. Simply labeling novel objects as ”Others” is often insufficient in safety-critical scenarios where object-specific recognition is necessary for decision-making.
Figure 1: GCFS in 3D object detection aims to adapt the source-pretrained model for strong performance on common and novel classes in the target domain via limited target samples.
Table 1: Summary of common 3D Object Detection Datasets, where the most common detection categories are highlighted by blue.
@@
KITTI [10] Karlsruhe (Germany) $7$ Car, Pedestrian, Truck, Van, Person_sitting, Cyclist, Tram
@@
NuScenes [1] Boston (USA), Singapore $23$ Car, Pedestrian, Truck, Bicycle, Barrier, Construction_vehicle, etc.
Waymo [36] 6 cities in USA $4$ Vehicle(including car, truck, and bus), Pedestrian, Cyclist, Sign
A2D2 [11] 50 cities in Germany $14$ Car, Pedestrian, Truck, Bicycle, Bus, UtilityVehicle, etc.
@@
To bridge the gap from dataset-based training to real-world deployment, we define a new task, generalized cross-domain few-shot learning (GCFS), for LiDAR-based 3D object detection. As conceptualized in Fig. 1, the GCFS task explicitly addresses both cross-domain adaptation and few-shot learning for novel categories, offering a cost-effective solution for rapid deployment in diverse environments. Unlike existing 3D few-shot learning (FSL)  [53, 37, 18], or its extension, 3D generalized few-shot learning (GFSL) [19], which assume the same distribution between training and deployment environments, GCFS accommodates both domain discrepancies and novel object adaptation while maintaining robust performance on common categories.
@@
Specifically, in the GCFS task, a 3D object detection model is initially trained with a source dataset that includes common object categories along with other possible source-specific classes. In the target deployment environment which may have domain gap from the source environment due to environmental factors, sensor configurations, and object appearances [48, 13, 17], we assume the presence of additional target-specific classes (i.e., novel classes) alongside the common classes. Given the practical feasibility and high cost of LiDAR data collection and annotation, we further assume that access to annotated data in the target environment is restricted to only a minimal amount (e.g., few-shot samples). The GCFS task, therefore, aims to enable the pre-trained model to adapt with minimal supervision in the target environment, ensuring strong performance on both common and target-specific novel categories. Though certain tasks in 2D object detection, like few-shot domain adaptation [9, 25] and generalized few-shot learning [7, 52], offer methodological insights into combining limited data adaptation with domain gap bridging, extending these 2D solutions effectively to the 3D domain remains challenging due to the higher-dimensional complexity and unique spatial characteristics of 3D data.
In this work, we introduce the first effective solution that integrates multi-modal fusion with contrastive-enhanced prototype learning to holistically overcome data scarcity and domain adaptation challenges in the GCFS task. Our multi-modal fusion module leverages vision-language models (VLMs) to extract open-set semantic information from point-cloud aligned images, enabling image-viewed 3D box proposals for improved object recall. However, accurate projection of 2D proposals into 3D space is challenged by point distribution biases that arise due to variations in object shapes and the inherent scanning nature of LiDAR sensors. To address this, we introduce a novel physically-aware 3D box searching strategy, which incorporates laser imaging principles to generate precise box proposals, improving localization across objects with diverse structural complexities. Though our image-guided fusion module significantly improves RPN’s proposal accuracy, challenges such as domain gaps, sparse point density, and limited data still hinder stable object feature learning through simple fine-tuning. To efficiently capture domain-specific representations for each object class in the target domain, we introduce contrastive-enhanced prototype learning, leveraging few-shot examples to encode object knowledge into flexible prototype embeddings. During inference, these prototypes act as anchors for feature refinement, adaptively aligning target-domain features with learned representations, thereby enhancing detection performance across domain shifts. In evaluation, we design three GCFS benchmark settings and conduct extensive experiments to illustrate the effectiveness of our solution. In sum, our contributions are as follows:
  • We formulate the first generalized cross-domain few-shot task in 3D object detection, comprehensively addressing the challenges in deploying dataset-pretrained models in real-world environments with diverse object categories.
  • We introduce the first GCFS solution for 3D object detection. It integrates multi-modal fusion with contrastive-enhanced prototype learning to overcome data scarcity and domain adaptation challenges holistically.
  • In our image-guided multi-modal fusion module, we propose a novel physical-aware box searching strategy for 2D-3D box proposal, boost the recall of objects.
  • We propose a novel contrastive-enhanced prototype learning strategy, enhancing model’s ability to swiftly adapt to the target domain with limited target data.
  • We establish three benchmark settings for evaluating GCFS tasks in 3D object detection, providing a comprehensive assessment framework for future exploration.
@@ 2.1 LiDAR-based 3D Object Detection
LiDAR-based 3D object detection [29, 23] aims to locate and classify objects of interest from input point clouds. Its models have mainly developed through point-based, voxel-based, and point-voxed-based. Point-based models [27, 31, 34] incorporate raw points and the PointNet-based backbones for fine-grained representation at the point level, albeit with high computational demands. Voxel-based methods [46, 22, 5, 54] represent the point cloud within a structured voxel grid and utilize sparse convolution for feature extraction, offering a trade-off between computational efficiency and spatial resolution. Point-voxed-based methods [32, 33] combine both, achieving a balance between efficiency and representation resolution, but often coming with increased model complexity and computation.
@@ 2.2 Few-shot Learning in Object Detection
In object detection, FSL aims to enable models to detect objects-of-interest with limited/few-shot labeled samples. In 2D tasks, extensive studies [44] address the challenge of data scarcity by leveraging techniques like meta-learning [45, 30], transfer learning [39, 3], and data augmentation [42]. In 3D object detection, most works focus on indoor scenarios. Based on VoteNet [28], Zhao et al. [53] introduce a prototypical vote module for local features refinement and a prototypical head module for global feature enhancement. Tang et al. [37] design a VAE-based prototype learning on top of it, and Li et al. [18] further leverage contrastive learning to learn more refined prototypical representations. However, extending 3D indoor object detection methods to outdoor scenarios is challenging due to sparse point clouds at greater distances, dynamic objects, and varying lighting and weather conditions. Recently, Liuet al. [19] propose the first outdoor generalized FSL solution for novel class learning. Yet without dealing with domain gaps in cross-domain scenarios, it leads to limited performance on GCFS settings.
@@ 2.3 Domain Adaptation in 3D Object Detection
The study of domain adaptation in 3D object detection mainly focus on unsupervised or semi-supervised settings. Corral-Soto et al. [4] integrate CycleGAN into the detection model, leveraging target-domain data with a small fraction of labeled frames to adapt the model to the target domain. Yang et al. [48, 3] employ a hybrid quality-aware triplet memory to generate pseudo-labels for unlabeled target-domain data. Hegde et al. [13] propose a source-free unsupervised DA approach that utilizes class prototypes to suppress noisy pseudo-labels on target data. These methods rely on a large amount of target data for domain adaptation and cannot deal with novel classes in the target domain, making them inapplicable to the GCFS task.
@@ 2.4 Open-vocabulary 3D Object Detection
Recently, open-vocabulary object detection [43, 49, 12, 50, 15] is garnering significant attention. In 3D object detection, these methods usually leverage 2D VLMs to acquire novel open-set semantics and enable detection models in identifying novel objects without annotations [21, 51, 6] . For instance, Lu et al. [21] propose to utilize CLIP-based VLMs to connect open-set textual knowledge and point-cloud representations for novel object identification. Najibi et al. [24] auto-label point cloud sequences using a pretrained 2D VLM and enable open-set training via vision-language knowledge distillation. Zhang et al. [51] introduce 2D-3D co-modeling that estimates corresponding 3D boxes from 2D insights with temporal and spatial constraints. Etchegaray et al. [6] employ 2D VLMs to generate pseudo-labels, augmented with simulation data. Since these 3D-OVD methods rely on large volumes of target data (including novel objects), their performance on the GSFS task remains to be validated.
3 Preliminaries: Generalized Cross-domain Few-Shot Learning
Figure 2: Proposed GCFS Framework. We first pretrain a detection model with source data. During model finetuning using target few-shot data, pairs of image and point cloud are processed by GDino+SAM and 3D backbone for image masks and 3D features. Insights from 2D analysis contribute in 1) enriching 3D features with 2D context and 2) proposing high-quality proposals via novel 2D-to-3D box search. Proposal features after RoI pooling are refined by learnable prototypes with attention mechanism, and then passed to final prediction.
Generalized Cross-domain Few-shot 3D object detection is a novel task designed to address the challenges of deploying 3D object detection models in real-world scenarios, where models trained on a source dataset must adapt to a new, target environment with limited labeled data across diverse categories. It is defined by two primary constraints. First, it assumes that only a limited number of labeled samples (i.e., few-shot samples) are available for all classes in the target environment, including both common classes (those shared with the source domain) and novel, target-specific classes. This reflects the practical limitations in collecting extensive data for every class in a new environment. Second, the GCFS setting accounts for a domain gap between the source and target environments, which may differ significantly in terms of object appearance, environmental conditions, and sensor setups. This gap presents an additional challenge for generalizing the model’s performance. Compared to FSL, DA, and OVD tasks, GCFS requires the solution to holistically address challenges of data scarcity and domain gap in real-world deployment of 3D object detection.
To formulate GCFS of 3D object detection, we distinguish LiDAR data from the source dataset and target environment (dataset) with superscript $s$ and $t$, respectively. In the source dataset used to pre-train a detection model $M_{pretrained}$, we assume sufficient amount of annotated data $D^{s}=\{\textbf{B}^{s}_{i},\textbf{C}^{s}_{i},\textbf{P}^{s}_{i}\mid\textbf{P}_{i}^{s}\in\mathbb{R}^{N\times 3}\}$ is available for the $\textbf{C}^{s}$ categories of objects-of-interest, where $\textbf{B}^{s}_{i}=\{\textbf{b}^{s}\mid\textbf{b}^{s}=[x,y,z,h,w,l,\theta]\}_{l=1}^{L_{i}}$ are the box annotations in point cloud $\textbf{P}_{i}^{s}$. In terms of target dataset $D^{t}=\{\textbf{B}^{t}_{i},\textbf{C}^{t}_{i},\textbf{P}^{t}_{i}\}$, only limited amount of data (e.g. few-shots) of the $\textbf{C}^{t}$ categories of object-of-interest is available. Here, we assume some categories in $\textbf{C}^{t}$ same as those in $\textbf{C}^{s}$, so certain knowledge in the pretrained model is valuable to the target task. In mathematics, these common classes are defined by $C_{com}=\textbf{C}^{t}\cap\textbf{C}^{s}\neq\emptyset$. We use $C_{nov}^{s}=\textbf{C}^{s}\setminus C_{com}$ and $C_{nov}^{t}=\textbf{C}^{t}\setminus C_{com}$ to denote the domain-specific classes, or novel classes. That is, objects in $C_{nov}^{t}$ are unseen in the source dataset. The target of the GCFS task is to obtain a strong detection model $M_{finetuned}$ through refining the model $M_{pretrained}$ with the $K$-shot examples in $D^{t}$.
@@ 4 Methodology
Fig. 2 illustrates the systematic diagram of our holistic solution for addressing the GCFS problem in LiDAR-based 3D object detection. It integrates multi-modal fusion (4.1) with contrastive-enhanced prototype learning (4.2) to overcome data scarcity and domain adaptation challenges. Briefly, our solution begins by pre-training a standard 3D object detection model on source data. In the target domain finetuning, a multi-modal fusion module leverages VLMs to extract rich, open-set semantic features from images. We specifically design a physically-aware box searching module to supplement the 3D region proposal network (RPN). Furthermore, to efficiently capture domain-specific representations for each class, we introduce contrastive-enhanced prototype learning supported by the few-shot examples. During inference, these prototypes serve as anchors for feature refinement, allowing the model to adaptively align target features with learned representations, thereby strengthening detection robustness and accuracy across domain gaps. This combination of multi-modal fusion and prototype-based feature refinement enables the model to generalize effectively to the target domain with minimal labeled data, addressing the unique requirements of the GCFS setting for real-world 3D object detection applications.
4.1 Image-guided Multi-modal Fusion
Trained on large-scale datasets, 2D Region VLMs [16, 14, 20] offer strong generalization capabilities to understand 2D visual information at a fine-grained, region-based level. In our GCFS task, we leverage point-cloud-aligned images and 2D VLMs, specifically, SAM [14] and Ground DINO (GDino) [20], to supplement the standard 3D detection. It not only enhances object detection (especially those in the novel classes) under the few-shot constraint, but also helps bridge the domain gap between source and target environments. In specifical, our image-guided multimodal fusion enhances the RPN’s object recall in twofolds. First, numerical features from the image analysis branch are integrated into the point-cloud representation, enriching it with 2D contextual information. Second, the 2D bounding boxes and masks are re-projected into 3D space to align with the point cloud data through our innovative physical-aware box searching. This re-projection addresses biases in point distributions across different object shapes that is often overlooked by existing box searching methods [6, 51]. By leveraging principles of laser imaging, our physical-aware 3D box searching strategy accurately generates 3D box proposals for objects of varying structural complexities.
Image-point feature fusion. In the point cloud branch, point cloud P is passed through the backbone encoder for point feature $\textbf{F}^{3D}\in\mathbb{R}^{N_{voxel}\times C}$. Simultaneously in the image branch, given the input image $\textbf{I}\in\mathbb{R}^{H\times W\times 3}$, we take object categories (including $C_{com}$ and $C_{nov}^{t}$) as text prompts to trigger GDino for detecting potential objects of interest with 2D boxes $\textbf{B}^{2D}\in\mathbb{R}^{N_{obj}\times 4}$. After the Non-maximum suppression (NMS) and score filtering, SAM takes $\textbf{B}^{2D}$ as box prompts to generate the object mask $\textbf{M}^{2D}\in\mathbb{R}^{H\times W\times|C^{t}|}$. Then, we utilize 2D-3D projection transformation on the coordinates of voxels $\textbf{P}^{voxel}$ to find corresponding 3D voxels of $\textbf{M}^{2D}$, obtaining voxel-wise feature $\textbf{F}^{voxel}\in\mathbb{R}^{N_{voxel}\times|C^{t}|}$,
$\textbf{F}^{voxel}=f_{proj}(\textbf{M}^{2D},\textbf{F}^{3D},\textbf{P}^{voxel}),$ (1)
where $f_{proj}$ denotes the re-projection function, defined by the intrinsic of the camera and extrinsic between the camera and LiDAR. To enrich the 3D features with 2D context information, we utilize a simple MLP to align channels of $\textbf{F}^{voxel}$ and $\textbf{F}^{3D}$ for a comprehensive feature $\textbf{F}^{fused}$,
$\textbf{F}^{fused}=\textbf{F}^{3D}+\textit{MLP}(\textbf{F}^{voxel})$ (2)
This feature fusion improve the following RPN for a better recall on objects of interest.
Physical-aware box searching. To find fine-grained box location from the 2D masks, we first take raw points as input to identify points within image masks,
$^{pts}=f_{proj}(\textbf{M}^{2D},\textbf{P}),$ P ) , (3)
where $\textbf{P}^{pts}$ is the point set corresponding to $\textbf{B}^{2D}$. We denote the point cloud of the $i^{th}$ object in $\textbf{B}^{2D}$ as $\textbf{P}_{i}\in\textbf{P}^{pts}$, and use the mean and 2×2$\times$2 ×-standard-deviation of all these point coordinates as the center and the boundary of the valid range to wipe out possible background points. For each class $c\in\textbf{C}^{t}$, we pre-define an anchor box $[h^{c},w^{c},l^{c}]$ according to mean object sizes of target few-shot data. The goal of box searching is to find the optimal location $[x,y,z]$ and heading angle $\theta$ for each $\textbf{P}_{i}$. Specifically, the variables $[x,y,z,\theta]$ shapes a reverse transformation T from $\textbf{P}_{i}$ to local coordinates $\textbf{P}^{local}_{i}$, $\textbf{P}^{local}_{i}=\textbf{T}\textbf{P}_{i}$, where
To find the right set of $[x,y,z,\theta]$, we design an outside distance loss $L_{OD}$ to ensure $\textbf{P}^{local}_{obj}$ are inside the box,
$L_{OD}=\sum_{c=1}^{C_{cls}}\sum_{\textbf{P}^{local}_{obj}\in c}min(|\textbf{P}^{local}_{i}|-\textbf{BD}^{c},0).$ (5)
Here, $\textbf{BD}^{c}=[h^{c}/2,w^{c}/2,l^{c}/2]$ denotes the boundary of local points from class $c$.

### v2

@@ Abstract
LiDAR-based 3D object detection models often struggle to generalize to real-world environments due to limited object diversity in existing datasets. To tackle it, we introduce the first generalized cross-domain few-shot (GCFS) task in 3D object detection, aiming to adapt a source-pretrained model to both common and novel classes in a new domain with only few-shot annotations. We propose a unified framework that learns stable target semantics under limited supervision by bridging 2D open-set semantics with 3D spatial reasoning. Specifically, an image-guided multi-modal fusion injects transferable 2D semantic cues into the 3D pipeline via vision-language models, while a physically-aware box search enhances 2D-to-3D alignment via LiDAR priors. To capture class-specific semantics from sparse data, we further introduce contrastive-enhanced prototype learning, which encodes few-shot instances into discriminative semantic anchors and stabilizes representation learning. Extensive experiments on GCFS benchmarks demonstrate the effectiveness and generality of our approach in realistic deployment settings.
Codehttps://github.com/Castiel-Lee/GCFS-3Det
@@ 1 Introduction
LiDAR-based 3D object detection (Zhang et al. 2025c; Baur et al. 2024; Mao et al. 2023) has significantly advanced autonomous driving by leveraging annotated datasets collected across diverse global locations (Geiger et al. 2012; Caesar et al. 2020; Sun et al. 2020; Geyer et al. 2020). However, as summarized in Table 1, existing datasets primarily focus on a limited set of common object categories (such as cars, pedestrians, and bicycles) within selected urban areas (e.g., USA, Singapore, and German cities). In contrast, real-world deployment introduces new geographic regions and novel object categories, such as electric scooters in Chinese cities or tuk-tuks in Thailand. Collecting and annotating large-scale LiDAR datasets for each new environment is both time-consuming and resource-prohibitive, which makes it unsuitable for rapid adaptation. This practical limitation highlights the need for methods that can generalize beyond the constraints of existing datasets: adapting to new domains and emerging object categories with minimal supervision.
Despite growing interest in these challenges, existing LiDAR-based 3D detection methods still face key limitations in effectively generalizing to novel categories with limited target-domain data. Among existing approaches, semi-supervised learning (Wang et al. 2023) and 3D open-vocabulary detection (OVD) (Etchegaray et al. 2024; Zhang et al. 2025a; Cao et al. 2024) often assume the availability of large amounts of unlabeled target data, which isn’t always feasible in model deployment. While 3D domain adaptation (DA)  (Wang et al. 2020b; Yang et al. 2022; Hegde and Patel 2024) focuses on addressing domain shifts, it does not explicitly account for novel object categories unseen during training. Simply labeling novel objects as ”others” is often insufficient in safety-critical scenarios where object-specific recognition is necessary for decision making.
Figure 1: GCFS in 3D object detection aims to adapt source-pretrained models for strong performance on common and novel classes in the target domain via limited target samples.
@@
KITTI (2012) Karlsruhe (Germany) 77 Car, Pedestrian, Truck, Van, Person_sitting, Cyclist, Tram
@@
NuScenes (2020) Boston (USA), Singapore $23$ Car, Pedestrian, Truck, Barrier, Construction_vehicle, etc.
Waymo (2020) 3 cities in USA 44 Vehicle(car, truck, and bus), Pedestrian, Cyclist, Sign
Argoverse 2 (2023) 6 cities in USA $30$ Car, Pedestrian, Truck, Bicycle, Motorcycle, Bus, Barrel, etc.
A2D2 (2020) 50 cities in Germany $14$ Car, Pedestrian, Truck, Bicycle, Bus, UtilityVehicle, etc.
Table 1: Summary of common 3D Object Detection Datasets, where the most common detection categories are underlined.
@@
To bridge the gap from dataset-based training to real-world deployment, we tackle a new task, generalized cross-domain few-shot (GCFS) learning, for LiDAR-based 3D object detection. As conceptualized in Fig. 1, the GCFS task comprehensively considers efficient adaptation to the target domain and stable semantic learning for novel and common categories via minimal target supervision, offering a cost-effective solution for rapid deployment in diverse environments. Unlike existing 3D few-shot learning (FSL)  (Zhao and Qi 2022; Tang et al. 2024; Li et al. 2024), or its extension, 3D generalized few-shot learning (GFSL) (Liu et al. 2023), which assumes the same distribution between training and deployment environments, GCFS accommodates both domain discrepancies and semantic adaptation target under limited target supervision.
@@
Specifically, in GCFS tasks, a 3D object detection model is initially trained on a source dataset including common object classes along with other possible source-specific classes. In the target environment, which may have a domain gap from the source data due to environmental factors, sensor configurations, and object appearances (Yang et al. 2022; Hegde and Patel 2024; Li et al. 2025), we assume the presence of additional target-specific classes (i.e., novel classes) alongside the common classes. Given the practical feasibility and high cost of LiDAR data collection and annotation, we further assume that access to annotated data in the target environment is restricted to only a minimal amount (e.g., few-shot samples). The GCFS task, therefore, aims to enable the pre-trained model to adapt with minimal supervision in the target environment, ensuring strong performance on both common and target-specific novel categories. Although certain tasks in 2D object detection, such as few-shot domain adaptation (Gao et al. 2023; Nakamura et al. 2022) and generalized few-shot learning (Fan et al. 2021; Zhang et al. 2023b), offer methodological insights into combining limited data adaptation with domain gap bridging, extending these 2D solutions effectively to the 3D domain remains challenging due to the higher-dimensional complexity and unique spatial characteristics of 3D data.
In this work, we introduce the first effective solution to comprehensively address the challenge of stable semantic representation learning under minimal target supervision in GCFS tasks. Our key insight is that generalization across domains and object categories is possible by bridging 2D open-set semantics and 3D spatial reasoning. By aligning sparse 3D observations with rich 2D vision-language priors and refining object understanding through prototype-based semantic anchoring, models can adapt robustly to both domain shifts and novel object classes from a few labeled examples. To realize this, we propose a unified GCFS framework built on two synergistic components: (1) an image-guided multi-modal fusion module that injects transferable 2D semantic cues into the 3D detection pipeline, improving proposal quality even in sparse point clouds; and (2) a contrastive-enhanced prototype learning mechanism that encodes few-shot target samples into discriminative, class-specific semantic anchors. Notably, we introduce a physically-aware box search strategy to improve 2D-to-3D alignment, and use contrastive learning to stabilize semantic prototypes under limited data. Together, these components enable robust adaptation with minimal supervision, offering a practical and generalizable solution for real-world 3D object detection. In evaluation, we design four GCFS benchmark settings and conduct extensive experiments to illustrate the effectiveness of our solution. In sum, our contributions are:
  • We formulate the generalized cross-domain few-shot task for 3D object detection and propose the first GCFS solution, holistically addressing domain shifts and novel object categories under limited supervision.
  • We propose a unified framework that leverages image-guided semantic grounding and contrastive prototype refinement to learn transferable object-level representations from sparse 3D data. Our framework illustrates that combining 2D vision-language priors with 3D geometry and few-shot semantic anchoring enables robust generalization across diverse environments and categories.
  • We establish four GCFS benchmark settings and show that our approach outperforms existing methods, providing a standardized framework for future research on 3D detection under domain and data constraints.
@@ 2.1 LiDAR-based 3D Object Detection
LiDAR-based 3D object detection (Zhang et al. 2025c; Gambashidze et al. 2024; Mao et al. 2023) aims to locate and classify objects of interest from input point clouds. Its models are primarily categorized into point-based, voxel-based, and point-voxel-based methods. Point-based models (Pan et al. 2021; Shi et al. 2019; Shi and Rajkumar 2020) incorporate raw points and the PointNet-based backbones for fine-grained representation at the point level, albeit with high computational demands. Voxel-based methods (Yan et al. 2018; Mao et al. 2021; Deng et al. 2021; Zhou and Tuzel 2018) represent the point cloud within a structured voxel grid and utilize sparse convolution for feature extraction, offering a trade-off between computational efficiency and spatial resolution. Point-voxel-based methods (Shi et al. 2023, 2020) combine both, achieving a balance between efficiency and representation resolution, but often coming with increased model complexity and computation.
@@ 2.2 Few-shot Learning in Object Detection
In object detection, FSL aims to enable models to detect objects with limited labeled samples. In 2D, extensive studies (Zhang et al. 2025b; Xin et al. 2024) tackle data scarcity by exploiting techniques like meta-learning (Yan et al. 2019; Ren et al. 2022), transfer learning (Wang et al. 2020a; Chen et al. 2018), and data augmentation (Wu et al. 2020). In 3D object detection, most works focus on indoor scenarios. Based on VoteNet (Qi et al. 2019), Proto-Vote (Zhao and Qi 2022) introduces a prototypical vote module for local features refinement and a prototypical head module for global feature enhancement. On top of it, a VAE-based prototype learning (Tang et al. 2024) is designed, and contrastive learning  (Li et al. 2024) is further exploited to learn more refined prototypical representations. However, extending 3D indoor object detection methods to outdoor scenarios is challenging due to sparse point clouds at greater distances, dynamic objects, and varying lighting and weather conditions. A recent work  (Liu et al. 2023) proposes the first outdoor generalized FSL solution for novel class learning. Yet, without dealing with domain gaps in cross-domain scenarios, it leads to limited performance on GCFS settings.
@@ 2.3 Domain Adaptation in 3D Object Detection
The study of domain adaptation in 3D object detection mainly focuses on unsupervised or semi-supervised settings. Works (Yang et al. 2022; Chen et al. 2018) employ a hybrid quality-aware triplet memory to generate pseudo-labels for unlabeled target-domain data. A source-free unsupervised DA approach (Hegde and Patel 2024) utilizes class prototypes to suppress noisy pseudo-labels on target data. Density-resampling-based augmentation and test-time adaptation (Li et al. 2025) are proposed to bridge density-related domain gaps. Yet, dependence on large target datasets and the inability to handle novel classes make these methods inapplicable to GCFS tasks
@@ 2.4 Open-vocabulary 3D Object Detection
Recently, open-vocabulary object detection (Wu et al. 2024; Zareian et al. 2021; Gu et al. 2021; Zhang et al. 2023a; Li* et al. 2022) has garnered significant attention. In 3D object detection, these methods usually take advantage of 2D VLMs to acquire novel open-set semantics and enable detection on novel objects without annotations. For instance, Lu et al. 2023 proposes to utilize CLIP-based VLMs to connect open-set textual knowledge and point-cloud representations for novel object identification. Auto-label methods  (Najibi et al. 2023; Etchegaray et al. 2024) are applied to point cloud sequences via a pretrained 2D VLM and enable novel semantic discovery for self-training. A 2D-3D co-modeling approach (Zhang et al. 2025a) estimates corresponding 3D boxes from 2D insights with temporal and spatial constraints. Since these 3D-OVD methods rely on large volumes of target data (including novel objects), their performance on the GCFS task remains to be validated.
Figure 2: Proposed GCFS Framework. We first pretrain a detection model with source data. During model finetuning using target few-shot samples, each query—the image and point cloud pair—is processed by GDino+SAM and 3D backbone for 2D instance-level masks and 3D features (top block). Insights from 2D context contribute to 1) enriching 3D features $F^{\text{fused}}$ with 2D semantic clues and 2) proposing high-quality “Box Candidates” via a novel 2D-to-3D box search. Proposal features $\textbf{F}^{\text{prp}}$ are refined by learnable prototypes $\textbf{F}^{\text{pro}}$ with an attention mechanism, and then passed to the final prediction (bottom block).
@@ 3 Methodology
Problem Statement: To formulate the GCFS of 3D object detection, we distinguish LiDAR data from the source dataset and target environment (dataset) with superscripts $s$ and $t$, respectively. In the source dataset used to pre-train the model $M_{\text{pretrained}}$, we assume access to sufficient annotated data $\mathbb{D}^{s}=\{\textbf{B}^{s}_{i},\textbf{C}^{s}_{i},\textbf{P}^{s}_{i}\}_{i=1}^{N^{s}}$, where $\textbf{P}^{s}_{i}\in\mathbb{R}^{N{\text{pts}}\times 3}$ denotes the point cloud, $\textbf{B}^{s}_{i}=\{\textbf{b}^{s}\mid\textbf{b}^{s}=[x,y,z,h,w,l,\theta]\}_{l=1}^{N_{\text{obj}}}$ the 3D bounding boxes, and $\textbf{C}^{s}_{i}$ the corresponding object category belonging to the source category space $\mathbb{C}^{s}$. For the target dataset $\mathbb{D}^{t}={(\textbf{B}^{t}_{i},\textbf{C}^{t}_{i},\textbf{P}^{t}_{i})}_{i=1}^{N^{t}}$, only limited (few-shot) samples are available for each target object category in the target category set $\mathbb{C}^{t}$. Here, we assume some categories are shared in $\mathbb{C}^{t}$ and $\mathbb{C}^{s}$, so certain knowledge in $M_{\text{pretrained}}$ is valuable to the target task. Formally, these common classes are defined by $\mathbb{C}_{\text{com}}=\mathbb{C}^{t}\cap\mathbb{C}^{s}\neq\emptyset$. We use $\mathbb{C}_{\text{nov}}^{s}=\mathbb{C}^{s}\setminus\mathbb{C}_{\text{com}}$ and $\mathbb{C}_{\text{nov}}^{t}=\mathbb{C}^{t}\setminus\mathbb{C}_{\text{com}}$ to denote the domain-specific novel classes. That is, objects belonging to $\mathbb{C}_{\text{nov}}^{t}$ are unseen in the source dataset. The goal of GCFS tasks is to obtain a strong detection model $M_{\text{finetuned}}$ through refining $M_{\text{pretrained}}$ with the $K$-shot examples in $\mathbb{D}^{t}$.
Fig. 2 presents an overview of our framework. To learn stable target semantics under limited supervision, we integrate two key components: an image-guided multi-modal fusion module and a class-specific contrastive prototype learning module. The fusion module exploits vision-language models (VLMs) to extract open-set semantic cues from point-cloud-aligned images, guided by a physically-aware box searching strategy that models LiDAR scanning behavior in the 3D geometric space. Meanwhile, the prototype learning module encodes class-level semantics from few-shot target samples into discriminative prototype anchors, which refine and align object features during inference.
3.1 Image-guided Multi-modal Fusion (IMMF)
In GCFS tasks, detectors trained on source data must adapt to new domains and categories via minimal target supervision. Yet, LiDAR data, which is sparse and geometry-focused, offers limited semantic richness, especially for novel objects. In contrast, aligned RGB images offer dense, transferable visual features and access to open-set semantics via pre-trained VLMs. To bridge this semantic gap, we introduce an image-guided multi-modal fusion that enriches 3D point representations with 2D semantic cues extracted from Grounding DINO (GDino) (Liu et al. 2024) and SAM (Kirillov et al. 2023), improving detection robustness under domain and category shifts.
Image-guided feature fusion. Given the point cloud P, we extract the non-empty voxel feature $\textbf{F}^{\text{voxel}}\in\mathbb{R}^{N_{\text{voxel}}\times C}$ via a 3D backbone, where $C$ denotes the 3D feature dimension. For the paired image $\textbf{I}\in\mathbb{R}^{H\times W\times 3}$, we use object category names (i.e., $\mathbb{C}_{\text{com}}$ and $\mathbb{C}_{\text{nov}}^{t}$) as text prompts to activate GDino, producing $N_{\text{obj}}$ 2D boxes $\textbf{B}^{\text{2D}}\in\mathbb{R}^{N_{\text{obj}}\times 4}$ with class labels as potential semantic clues. After non-maximum suppression and confidence filtering, SAM takes $\textbf{B}^{\text{2D}}$ as box prompts and generates dense object masks $\textbf{M}^{\text{2D}}\in\mathbb{R}^{H\times W\times|\mathbb{C}^{t}|}$. We then project the coordinates $\textbf{P}^{\text{voxel}}$ of $\textbf{F}^{\text{voxel}}$ onto the image to identify the object masks and obtain the voxel-aligned object mask $\textbf{M}^{\text{vxl-obj}}\in\mathbb{R}^{N_{\text{voxel}}\times|\mathbb{C}^{t}|}$:
$\textbf{M}^{\text{vxl-obj}}=f_{\text{proj}}(\textbf{M}^{\text{2D}},\textbf{P}^{\text{voxel}}),$ (1)
where $f_{\text{proj}}(\cdot)$ denotes 2D-to-3D mapping based on known camera intrinsics and extrinsics. To integrate 2D semantic cues into the 3D representation, we apply an MLP to align the channel dimensions and fuse the features:
$\textbf{F}^{\text{fused}}=\textbf{F}^{\text{voxel}}+\text{MLP}(\textbf{M}^{\text{vxl-obj}}).$ (2)
This fused feature $\textbf{F}^{\text{fused}}$ enhances the downstream region proposal network (RPN), improving object recall for both common and novel categories.